OpenStudy (anonymous):

Find a polynomial f(x) of degree 3 that has the following zeros. 8,-1,0 Leave your answer in factored form.

OpenStudy (anonymous):

(x-8)(x+1)(x) = 0 now do the multiplication to get polynomial

OpenStudy (johnweldon1993):

The easy thing about leaving it in factored form is that you can just plug in the 3 numbers you have... your polynomial breaks down into (x - 8)(x + 1)(x) = 0 Because as you can see...you can plug in...8....or -1...or 0....and that entire equation will still = 0 Since this question states to leave it in factored form...you are done here

OpenStudy (anonymous):

f(x)=x(x-8)(x+1)

OpenStudy (anonymous):

x^3-7x^2-8x=0 Would I get this?